OpenStudy (anonymous):

two fifths of the 100 passengers stayed in the subway cars until the last stop.how many of the 100 passengers got off the subway cars before the last stop

OpenStudy (dumbcow):

if 2/5 did not get off then 3/5 must have gotten off 2/5 + 3/5 = 1 now find 3/5 of 100 \[\frac{3}{5}*\frac{100}{1}\] =300/5 =60

OpenStudy (anonymous):

its 60 or 50? because i got 50

OpenStudy (dumbcow):

should be 60 it cant be 50 because that would be one half but we know its 3/5

OpenStudy (anonymous):

oh ok i get it